# Algorithmic Aspects of Strategic Trading

### v1

@@ Abstract
Algorithmic trading in modern financial markets is widely acknowledged to exhibit strategic, game-theoretic behaviors whose complexity can be difficult to model. A recent series of papers (Chriss, 2024b, c, a, 2025) has made progress in the setting of trading for position building. Here parties wish to buy or sell a fixed number of shares in a fixed time period in the presence of both temporary and permanent market impact, resulting in exponentially large strategy spaces. While these papers primarily consider the existence and structural properties of equilibrium strategies, in this work we focus on the algorithmic aspects of the proposed model. We give an efficient algorithm for computing best responses, and show that while the temporary impact only setting yields a potential game, best response dynamics do not generally converge for the general setting, for which no fast algorithm for (Nash) equilibrium computation is known. This leads us to consider the broader notion of Coarse Correlated Equilibria (CCE), which we show can be computed efficiently via an implementation of Follow the Perturbed Leader (FTPL). We illustrate the model and our results with an experimental investigation, where FTPL exhibits interesting behavior in different regimes of the relative weighting between temporary and permanent market impact.
@@ 1 Introduction
In a series of recent papers, Chriss (Chriss, 2024b, c, a, 2025) has made significant progress with the introduction and analysis of a stylized but realistic model for competitive trading. In Chriss’ model, multiple traders play a game in which each player wishes to acquire either a long or short position in a common asset in a fixed time window, and each wishes to minimize their cost in doing so. As is standard in the finance literature, costs decompose into temporary and permanent market impact. Broadly speaking, temporary impact models the “mechanical” influence on prices inherent in the continuous double auction common in modern stock exchanges (due to worse prices as one eats further into the limit order books; see (Nevmyvaka et al., 2006) for background), while permanent impact models the longer-term “psychological” effects of trading, such as perceptions of the value of the underlying asset. As we shall see in the model, at equilibrium, temporary impact tends to make traders want to avoid each other temporally, while permanent impact tends to make traders want to “front run” (trade before) each other. These competing forces, along with the fact that it may be beneficial to sell some shares en route to acquiring a net long position, make for a rich set of equilibrium strategies, which is the primary focus of Chriss’ papers.
In this work, we focus on the algorithmic and learning aspects of Chriss’ model, and in doing so broaden the class of equilibria considered (specifically to coarse correlated equilibria (CCE), the class to which no-regret dynamics are known to converge). More precisely, our results include the following:
@@
  • Given that the general game seems to admit no special form, and thus computing Nash equilibria may be intractable, we next turn attention to computing (approximate) coarse correlated equilibria via no-regret dyanmics. We prove that despite the exponentially large strategy space, there is a computationally efficient implementation of Follow the Perturbed Leader (FTPL) for our game. In addition to their computational tractability, CCE are interesting in our setting due to the possibility of higher social welfare, and the suggestion that lightweight correlation at the exchanges themselves might render trading less costly.
@@ 1.1 Related Work
Key to Chriss’ model are standard notions of (temporary and permanent) market impact, on which there is a large literature; see (J.Gatheral, 2010; Gatheral, 2010; Hautsch and Huang, 2012; Zarinelli et al., 2015; Bouchaud et al., 2002; Webster, 2023) for a representative but very partial sample. Broadly speaking, this literature considers various models for how trading activity influences asset prices, implications of those models for trading strategies, and empirical validation. There is also a smaller body of work considering the algorithmic aspects of optimal trading, including machine learning approaches (Even-Dar et al., 2006; Ganchev et al., 2010; Nevmyvaka et al., 2006; Kakade et al., 2004). Our work is also focused on algorithmic considerations, but in a game-theoretic setting.
@@ 2 Model and Preliminaries
A trading strategy $a:[T]\to\mathbb{Z}$ is a mapping from a time step $t$ to the number of shares held by a player at time $t$, satisfying $a(0)=0$ and $a(T)=V$ for a target volume $V$.
@@
The cost of a trading strategy $a_{i}\in\mathcal{A}_{i}$ against strategies $a_{-i}\in\mathcal{A}_{-i}$ is given by:
@@
where $\kappa\geq 0$ is a market impact coefficient.
@@
We will study several basic equilibria concepts.
@@
An action profile $\bm{a}$ is an $\varepsilon$-approximate pure Nash equilibrium if for all $i$, $c(a_{i},a_{-i})\leq\min_{a\in\mathcal{A}_{i}}c(a,a_{-i})+\varepsilon$. When $\varepsilon=0$, we say $\bm{a}$ is a pure Nash equilibrium.
@@
We will also consider coarse correlated equilibria (CCE), a more general class of equilibria.
Definition 2.8 (Coarse Correlated Equilibrium).
@@ 3 Computing a Best Response
This idea lends itself to more general cost minimization problems over trading strategies; in fact, we will present and analyze a more general form of the algorithm than what is required to compute a best response in the trading game. In particular, we show how to compute:
$a^{*}=\operatorname*{arg\,min}_{a\in\mathcal{A}_{i}}\phi^{T}(a)$
for any cost function $\phi^{T}(a)$ that can be written as the sum of one-step costs $p^{t}$ that depend on the “state” of trades at time $t$:
@@
Given a target volume $V$, trading limits $\theta_{L},\theta_{U}$, and a cost function $\phi^{T}$ such that $\phi^{T}(a)=\sum_{t=1}^{T}p^{t}(a(t-1),a^{\prime}(t))$ for some function $p^{t}$, Algorithm 1 computes
$a^{*}=\operatorname*{arg\,min}_{a\in\mathcal{A}(V,\theta_{L},\theta_{U})}\phi^{T}(a)$
Moreover, Algorithm 1 runs in time $O((\theta_{U}-\theta_{L})^{2}T^{2})$.
@@
Why? The one-step cost of buying $k$ shares at time $t$ with $s$ shares remaining is $p^{t}(V-s,k)$; the minimum remaining cost is simply the minimum cost of buying $s-k$ shares beginning at the next time step, i.e. $\mathrm{OPT}(t+1,s-k)$.
The cost of a best response strategy is then $\mathrm{OPT}(1,V)$; some simple bookkeeping will allow us to recover the optimal strategy. We present these ideas in more detail below.
@@ 4 A Decomposition of the Trading Game
This decomposition shows that the basic structure of the trading game differs with underlying market impact. Most saliently, when $\kappa=0$ (i.e. when players face only temporary impact), the game is a potential game, and so simple dynamics converge to a pure Nash equilibrium. When $\kappa=2$ (i.e. the contribution of permanent impact is twice that of temporary impact), the game is constant-sum. For all other $\kappa$, the game is a weighted mixture of a potential game and a constant-sum game; the value of $\kappa$ determines its proximity to either. Later on, we will see that this basic structure is reflected in interesing ways in our experimental evaluations.
@@ 4.1 The Temporary Impact Regime
Consider an $n$-player game $G$ where each player $i\in[n]$ chooses actions from an action set $\mathcal{A}_{i}$. $G$ is an exact potential game if there exists a potential function $\phi:\mathcal{A}_{1}\times...\times\mathcal{A}_{n}\to\mathbb{R}$ such that for all players $i\in[n]$ and actions $a_{i},b_{i}\in\mathcal{A}_{i}$,
@@
We have thus established that a pure Nash equilibrium exists in the temporary impact only setting and can be found using best response dynamics — but how quickly? To answer this, it is common to settle for an $\varepsilon$-approximate equilibrium. We can bound the number of rounds best response dynamics will run for, as long as each deviation leads to a large enough improve in cost—at least $\varepsilon$. That is, as along as the strategy profile is not an $\varepsilon$-approximate Nash equilibrium, players sequentially move to a strategy that lowers their cost by at least $\varepsilon$. Notice we can implement best response dynamics using Algorithm 1 to sequentially compute best responses, and verifying if it gives a large enough improvement.
@@
Algorithm 2 returns an $\varepsilon$-approximate Nash equilibrium in the temporary impact only setting. Moreover, it has running time bounded by $O\left(\frac{n^{3}\theta^{4}T^{3}}{\varepsilon}\right)$, where $\theta=\max\{|\theta_{L}|,|\theta_{U}|\}$.
@@ 4.2 The Permanent Impact Regime
to be the cost that averages the permanent impact contribution from the previous time step and the current time step.
@@ 5 Efficient Equilibria Computation in the General Game
In the previous section, we established that the general trading game admits a meaningful decomposition; however, this has no immediate implications for tractable equilibria computation. In this section, we relax our goal to the computation of coarse correlated equilibria, a broader class of equilibria that encapsulates Nash equilibria. We show how to efficiently implement no-regret dynamics, for which the empirical history of play converges to a joint distribution that is an (approximate) coarse correlated equilibrium.
In no-regret dynamics, each player chooses a sequence of (randomized) actions by running a regret-minimizing algorithm over $R$ rounds.
@@
The (per-round) regret of a player $i$ who chooses a sequence of strategies $a_{i,1},...,a_{i,R}\in\mathcal{A}_{i}$ is defined as:
$Reg_{i}^{R}=\max_{a_{i}\in\mathcal{A}_{i}}\frac{1}{R}\sum_{r=1}^{R}(c(a_{i,r},a_{-i,r})-c(a_{i},a_{-i,r}))$
@@
No regret encodes the idea that if a player looks back at the history of play, they would find that no deviation to a fixed strategy would have improved their cost. It thus follows that if each player has regret bounded by $\varepsilon$, the time-averaged, empirical distribution of play is an $\varepsilon$-approximate coarse correlated equilibrium.
The difficulty in our setting is producing a no-regret algorithm that is computationally efficient, given that the space of trading strategies is exponentially large. A classic no-regret algorithm is the Multiplicative Weights algorithm (see (Arora et al., 2012) for an extensive survey), which achieves regret decreasing at a rate of $O(\sqrt{\ln|\mathcal{A}|/R})$. Thus, if all players run their own Multiplicative Weights algorithm, the empirical distribution of play converges to an $\varepsilon$-approximate coarse correlated equilibrium after $O(\ln|\mathcal{A}|/\varepsilon^{2})$ rounds. However, running Multiplicative Weights will be computationally expensive for our problem: it maintains an explicit distribution over the strategy space and so requires per-round computation that is linear in the number of strategies—which, for our setting, is exponential in $T$.
We show that an instantiation of the Follow The Perturbed Leader (FTPL) algorithm can be used to implement no-regret dynamics efficiently — the number of rounds required to reach an approximate equilibrium and the per-round runtime will be polynomial in our problem parameters. Next we introduce FTPL and state its regret guarantees.
@@
$\max_{f\in\mathcal{F}}\frac{1}{R}\sum_{r=1}^{R}\left(\mathbb{E}[\langle f_{r}),h_{r})\rangle]-\langle f,h_{r})\rangle\right)\leq 2\sqrt{\frac{DMC}{R}}$
@@
In order to implement FTPL, we must first write our problem as an instance of OLO. We will take the perspective of player $i$ (who corresponds to the learner) computing a no regret strategy against other players (who, in aggregate, correspond to the adversary). We “linearize” the problem by constructing low-dimensional strategy spaces for the learner and adversary as follows. The learner will play over the space $\mathcal{F}_{\mathcal{A}_{i}}=\{f(a_{i})\}_{a_{i}\in\mathcal{A}_{i}}\subseteq\mathbb{R}^{2T}$ and the adversary will play over the space $\mathcal{H}_{\mathcal{A}_{-i}}=\{h(a_{-i})\}_{a_{-i}\in\mathcal{A}_{-i}}\subseteq\mathbb{R}^{2T}$, where $h$ and $f$ apply the following transformations:
@@
We can see that the learner’s cost of playing $f(a_{i})$ against $h(a_{-i})$ in the OLO problem matches their cost of playing $a_{i}$ against $a_{-i}$ in the trading game:

### v2

@@ Abstract
Algorithmic trading in modern financial markets is widely acknowledged to exhibit strategic, game-theoretic behaviors whose complexity can be difficult to model. A recent series of papers (Chriss, 2024b, c, a, 2025) has made progress in the setting of trading for position building. Here parties wish to buy or sell a fixed number of shares in a fixed time period in the presence of both temporary and permanent market impact, resulting in exponentially large strategy spaces. While these papers primarily consider the existence and structural properties of equilibrium strategies, in this work we focus on the algorithmic aspects of the proposed model. We give an efficient algorithm for computing best responses, and show that while the temporary impact only setting yields a potential game, best response dynamics do not generally converge for the general setting, for which no fast algorithm for (Nash) equilibrium computation is known. This leads us to consider the broader notion of Coarse Correlated Equilibria (CCE), which we show can be computed efficiently via an implementation of Follow the Perturbed Leader (FTPL). While we focus on equilibrium computation, our FTPL implementation learns no-regret strategies in any (adversarial) trading environment. We illustrate the model and our results with an experimental investigation, where FTPL exhibits interesting behavior in different regimes of the relative weighting between temporary and permanent market impact.
@@ 1 Introduction
In a series of recent papers, Chriss (Chriss, 2024b, c, a, 2025) has made significant progress with the introduction and analysis of a stylized but realistic model for competitive trading. In Chriss’ model, multiple traders play a game in which each player wishes to acquire either a long or short position in a common asset in a fixed time window, and each wishes to minimize their cost in doing so. As is standard in the finance literature, costs decompose into temporary and permanent market impact. Broadly speaking, temporary impact models the “mechanical” influence on prices inherent in the continuous double auction common in modern stock exchanges (due to worse prices as one eats further into the limit order books; see (Nevmyvaka et al., 2006) for background), while permanent impact models the longer-term “psychological” effects of trading, such as perceptions of the value of the underlying asset. As we shall see in the model, at equilibrium, temporary impact tends to make traders want to avoid each other temporally, while permanent impact tends to make traders want to “front run” (trade before) each other. These competing forces, along with the fact that it may be beneficial to sell some shares en route to acquiring a net long position, make for a rich set of equilibrium strategies, which is the primary focus of Chriss’ papers [^2] Chriss’ work is closely related to Carlin et al. (2007), who study equilibria in similar competitive trading setup..
In this work, we focus on the algorithmic and learning aspects of Chriss’ model, and in doing so broaden the class of equilibria considered (specifically to coarse correlated equilibria (CCE), the class to which no-regret dynamics are known to converge). Our results include the following:
@@
  • Given that the general game seems to admit no special form, and thus computing Nash equilibria may be intractable, we next turn attention to computing (approximate) coarse correlated equilibria via no-regret dyanmics. We prove that despite the exponentially large strategy space, there is a computationally efficient implementation of Follow the Perturbed Leader (FTPL) for our game. In addition to their computational tractability, CCE are interesting in our setting due to the possibility of higher social welfare, and the suggestion that lightweight correlation at the exchanges themselves might render trading less costly. Beyond equilibrium computation, our implementation of FTPL can be used to learn no-regret strategies in any environment, where other traders could be acting adversarially.
@@ 1.1 Related Work
Key to Chriss’ model are standard notions of (temporary and permanent) market impact, on which there is a large literature; see (J.Gatheral, 2010; Gatheral, 2010; Hautsch and Huang, 2012; Zarinelli et al., 2015; Bouchaud et al., 2002; Webster, 2023) for a representative but partial sample. Broadly speaking, this literature considers various models for how trading activity influences asset prices, implications of those models for trading strategies, and empirical validation. A smaller body of work considers the algorithmic aspects of optimal trading, including machine learning approaches (Even-Dar et al., 2006; Ganchev et al., 2010; Nevmyvaka et al., 2006; Kakade et al., 2004). Our work also focuses on algorithmic considerations, but in a game-theoretic setting.
Several other works consider trading games but operate in models different from ours. Carlin et al. (2007) model players as either “buyers” who wish to buy or sell a predetermined number of shares over a fixed time period or “predators” seeking a zero net position, and they derive a pure Nash equilibrium. Similar to Chriss and our model, they use a linear model of temporary and permanent impact (with additional noise governed by a stochastic process). Cont et al. (2025) study a sequential-move game, deriving Stackelberg equilibria when a resource-rich but less informed institutional trader “leads,” and an information-rich but limited-resource high frequency trader “follows.” Like Chriss, both Carlin et al. (2007) and Cont et al. (2025) work in a continuous time and volume model.
@@ 2 Model and Preliminaries
A trading strategy $a:[T]\to\mathbb{Z}$ is a mapping from a time step $t$ to the number of shares held by a player at time $t$, satisfying $a(0)=0$ and $a(T)=V$.
@@
Fix $\kappa\geq 0$. The cost of a trading strategy $a_{i}\in\mathcal{A}_{i}$ against strategies $a_{-i}\in\mathcal{A}_{-i}$ is given by:
@@
It is worth noting that this model can be “explained” in terms of assumptions on the underlying limit order book dynamics that mediate all trading activity. More specifically, considering (without loss of generality) only players who wish to buy shares to obtain a long position, if we assume that (a) the distribution of share prices in the sell order book is uniform, and that (b) once consumed, shares in the sell book are never replenished by the arrival of new shares, then we recover Chriss’ model. Assumption (a) corresponds to his linear temporary cost model, and assumption (b) corresponds to his linear permanent cost model. Assumption (b) can be viewed as an extreme form of permanent impact, in that any trading activity that drives the price up will never be reversed — the market always revalues the security at the current price level. It is then possible to interpret $\kappa$ as a liquidity replenishment parameter, in that intermediate values of $\kappa$ model new sell orders arriving at previous price levels at some rate. Carlin et al. (2007) attribute permanent impact to a similar phenomenon—that decreases in supply cause increases in price—but do not directly map it onto limit order book dynamics.
While both of these assumptions are stylized and somewhat unrealistic in practice, they at least ground our model in assumptions about the low-level dynamics of the exchanges. They also point to more realistic variants of the model, in which we assume more natural price distributions in the order books (for instance, it is common for much more liquidity to be aggregated near the bid and ask prices, and to thin out away from them), and less extreme replenishment assumptions.
@@
We will study several basic equilibria concepts, defined below in increasing generality.
@@
An action profile $\bm{a}$ is an $\varepsilon$-approximate pure Nash equilibrium if for all $i$, $c(a_{i},a_{-i})\leq\min_{a\in\mathcal{A}_{i}}c(a,a_{-i})+\varepsilon$. When $\varepsilon=0$, $\bm{a}$ is a pure Nash equilibrium.
@@
Definition 2.8 (Correlated Equilibrium (CE)).
A distribution $\mathbf{D}$ over action profiles is an $\varepsilon$-approximate correlated equilibrium if for all $i$, for all swap functions $\phi_{i}:\mathcal{A}_{i}\to\mathcal{A}_{i}$, $\mathbb{E}_{\bm{a}\sim\mathbf{D}}[c(a_{i},a_{-i})]\leq\mathbb{E}_{\bm{a}\sim\mathbf{D}}[c(\phi(a_{i}),a_{-i})]+\varepsilon$. When $\varepsilon=0$, $\mathbf{D}$ is a correlated equilibrium.
Definition 2.9 (Coarse Correlated Equilibrium (CCE)).
@@ 3 Computing a Best Response
This idea lends itself to more general cost minimization problems over trading strategies; in fact, we will present and analyze a more general form of the algorithm than what is required to compute a best response in the trading game. In particular, we show how to compute $a^{*}=\operatorname*{arg\,min}_{a\in\mathcal{A}_{i}}\phi^{T}(a)$ for any cost function $\phi^{T}(a)$ that can be written as the sum of one-step costs $p^{t}$ that depend on the “state” of trades at time $t$:
@@
Given a target volume $V$, trading limits $\theta_{L},\theta_{U}$, and a cost function $\phi^{T}$ such that $\phi^{T}(a)=\sum_{t=1}^{T}p^{t}(a(t-1),a^{\prime}(t))$ for some function $p^{t}$, Algorithm 1 computes $a^{*}=\operatorname*{arg\,min}_{a\in\mathcal{A}(V,\theta_{L},\theta_{U})}\phi^{T}(a)$. Moreover, Algorithm 1 runs in time $O((\theta_{U}-\theta_{L})^{2}T^{2})$.
@@
Why? The one-step cost of buying $k$ shares at time $t$ with $s$ shares remaining is $p^{t}(V-s,k)$; the minimum remaining cost is simply the minimum cost of buying $s-k$ shares beginning at the next time step, i.e. $\mathrm{OPT}(t+1,s-k)$. The cost of a best response strategy is then $\mathrm{OPT}(1,V)$; some simple bookkeeping will allow us to recover the optimal strategy.
We present these ideas in more detail below.
@@ 4 A Decomposition of the Trading Game
This decomposition shows that the basic structure of the trading game differs with underlying market impact [^4] Empirical studies suggest that the ratio of temporary to permanent impact varies significantly in markets (Sadka, 2006; Carlin et al., 2007); our result highlights how incentives can vary with this ratio.. Most saliently, when $\kappa=0$ (i.e. when players face only temporary impact), the game is a potential game, and so simple “best-response dynamics” converge to a pure Nash equilibrium. When $\kappa=2$ (i.e. the contribution of permanent impact is twice that of temporary impact), the game is constant-sum. In this case, the empirical history of “no-regret dynamics” converges to a mixed Nash equilibrium (Cesa-Bianchi and Lugosi, 2006). For all other $\kappa$, the game is a weighted mixture of a potential game and a constant-sum game; the value of $\kappa$ determines its proximity to either. Later on, we will see that this basic structure is reflected in interesting ways in our experimental evaluations.
@@ 4.1 The Temporary Impact Regime
Consider an $n$-player game $G$ where each player $i\in[n]$ chooses actions from an action set $\mathcal{A}_{i}$ and has cost function $c_{i}$. $G$ is an exact potential game if there exists a potential function $\phi:\mathcal{A}_{1}\times...\times\mathcal{A}_{n}\to\mathbb{R}$ such that for all players $i\in[n]$ and actions $a_{i},b_{i}\in\mathcal{A}_{i}$,
@@
We have thus established that a pure Nash equilibrium exists in the temporary impact only setting and can be found using best response dynamics — but how quickly? To answer this, it is common to settle for an $\varepsilon$-approximate equilibrium. We can bound the number of rounds best response dynamics will run for, as long as each deviation leads to a large enough improve in cost—at least $\varepsilon$. That is, as long as the strategy profile is not an $\varepsilon$-approximate Nash equilibrium, players sequentially move to a strategy that lowers their cost by at least $\varepsilon$. Notice we can implement best response dynamics using Algorithm 1 to sequentially compute best responses, and verifying if it gives a large enough improvement.
@@
Best response dynamics (Algorithm 2) returns an $\varepsilon$-approximate Nash equilibrium in the temporary impact only setting. Moreover, it has running time bounded by $O\left(\frac{n^{3}\theta^{4}T^{3}}{\varepsilon}\right)$, where $\theta=\max\{|\theta_{L}|,|\theta_{U}|\}$.
@@ 4.2 The Permanent Impact Regime
to be the cost averaging the permanent impact contribution from the previous and current time step.
@@ 5 Efficient Equilibria Computation in the General Game
We have just established that the general trading game admits a meaningful decomposition; however, this has no immediate implications for tractable equilibria computation. In this section, we relax our goal to the computation of CCE, a broader class of equilibria that encapsulates Nash equilibria. We show how to efficiently implement no-regret dynamics, for which the empirical history of play converges to a joint distribution that is an (approximate) CCE. In no-regret dynamics, each player chooses a sequence of (randomized) actions by running a regret-minimizing algorithm over $R$ rounds.
@@
The (average) regret of a player $i$ who chooses a sequence of strategies $a_{i,1},...,a_{i,R}\in\mathcal{A}_{i}$ is defined as:
$Reg_{i}(R)=\max_{a_{i}\in\mathcal{A}_{i}}\frac{1}{R}\sum_{r=1}^{R}(c(a_{i,r},a_{-i,r})-c(a_{i},a_{-i,r}))$
@@
No regret encodes the idea that if a player looks back at the history of play, they would find that no deviation to a fixed strategy would have improved their cost. It thus follows that if each player has regret bounded by $\varepsilon$, the time-averaged, empirical distribution of play is an $\varepsilon$-approximate CCE.
The difficulty in our setting is producing a no-regret algorithm that is computationally efficient, given that the space of trading strategies is exponentially large. A classic no-regret algorithm is the Multiplicative Weights algorithm (see (Arora et al., 2012) for an extensive survey), which achieves regret decreasing at a rate of $O(\sqrt{\ln|\mathcal{A}|/R})$. Thus, if all players run their own Multiplicative Weights algorithm, the empirical distribution of play converges to an $\varepsilon$-approximate CCE after $O(\ln|\mathcal{A}|/\varepsilon^{2})$ rounds. However, running Multiplicative Weights will be computationally expensive for our problem: it maintains an explicit distribution over the strategy space and so requires per-round computation that is linear in the number of strategies—which, for our setting, is exponential in $T$.
We show that an instantiation of the Follow The Perturbed Leader (FTPL) algorithm (Kalai and Vempala, 2005) can be used to implement no-regret dynamics efficiently—the number of rounds required to reach an approximate equilibrium and the per-round runtime is polynomial in our problem parameters. We note that FTPL maintains no-regret guarantees in adversarial environments, so while our main focus is on the joint execution of FTPL by all players, our instantiation of FTPL can be used by a single player to obtain vanishing regret in any environment, where other players could be acting adversarially. Next we introduce FTPL and state its regret guarantees.
@@
$\max_{f\in\mathcal{F}}\frac{1}{R}\sum_{r=1}^{R}\left(\mathbb{E}[\langle f_{r},h_{r}\rangle]-\langle f,h_{r})\rangle\right)\leq 2\sqrt{\frac{DMC}{R}}$
@@
In order to implement FTPL, we must first cast our problem as an instance of OLO. We will take the perspective of player $i$ (who corresponds to the learner) computing a no regret strategy against other players (who, in aggregate, correspond to the adversary). Although players’ cost functions are not linear in their actions, we show how to “linearize” the problem by constructing higher-dimensional representations of the strategy spaces for the learner and adversary. Broadly speaking, we use the fact that the cost function is linear in the opponents’ strategies $a_{-i}$ and introduce dimensions to represent nonlinearities in $a_{i}$—in particular the product relationships. More specifically, the learner will play over the space $\mathcal{F}_{\mathcal{A}_{i}}=\{f(a_{i})\}_{a_{i}\in\mathcal{A}_{i}}\subseteq\mathbb{R}^{2T}$ and the adversary will play over the space $\mathcal{H}_{\mathcal{A}_{-i}}=\{h(a_{-i})\}_{a_{-i}\in\mathcal{A}_{-i}}\subseteq\mathbb{R}^{2T}$, where $h$ and $f$ apply the following transformations:
@@
Note that the dimension of the strategy space is only larger by a factor of 2. The transformation is cost-preserving: the learner’s cost of playing $f(a_{i})$ against $h(a_{-i})$ in the OLO problem matches their cost of playing $a_{i}$ against $a_{-i}$ in the trading game:
@@
In Appendix A, we investigate computation of approximate CE via no-swap-regret dynamics.